OpenStudy (anonymous):

The length of a rectangle is 6 yards longer than its width. If the perimeter of the rectangle is 32 yd, find its area in square yards.

OpenStudy (anonymous):

Area of a rectangle is L x W L - length W-width Perimeter is 2L + 2W Now they told us length is 6 yds longer than the width so L = W + 6 Now since the perimeter is 32 We know 2L + 2W = 32 Substituting L = W + 6 2(W+ 6) + 2W = 32 Distributing 2 W + 12 + 2W = 32 Combining like terms 4W + 12 = 32 Subtract 12 from both sides 4W = 20 Dividing both sides be 4 W = 5 Now we have W then L = W + 6 = 5 + 6 = 11 So L = 11 W = 5 so Area = L x W = 11 x 5 = 55 yd^2